# soft open fences

### v1

@@ 4 Discussion and Future Work
Additionally, the connection between entropy minimization and implicit confidence calibration warrants deeper investigation. Our findings suggest that EM enhances model confidence by reinforcing high-probability reasoning paths. This implies EM might serve as a lightweight alternative to complex calibration techniques, especially for tasks where interpretability and robustness are critical. Developing evaluation protocols to more precisely quantify EM’s calibration effects will be an important direction.
@@
Entropy minimization is conceptually orthogonal to most existing post-training paradigms, including supervised fine-tuning (SFT), reinforcement learning with human feedback (RLHF), and preference-based optimization. This opens exciting opportunities for hybrid methods. For example, EM can be applied before SFT to sharpen a model’s predictive distributions, improving its receptiveness to downstream supervision. Alternatively, EM could serve as a regularization strategy during SFT or RLHF, promoting confident but calibrated behavior without overfitting to reward signals.
@@ 5.1 Entropy Minimization
Agarwal et al.[16] was the first to discover that simply optimizing the entropy loss can lead to significant improvements in reasoning performance. However, they did not explore this phenomenon in depth, stating in the original paper, "We leave the rigorous analysis to future works." Wang et al.[1] is the first work to study entropy minimization in large language models. However, they consider the effectiveness of entropy minimization to be unreasonable, presenting experimental results but offering limited in-depth analysis of entropy minimization.
@@ 7 Acknowledgement
We thank Tian Xie, Zhengmao Ye, phwu and jdli at Ubiquant for their support and insightful discussions.

### v2

@@ 4 Future Work
Additionally, the connection between entropy minimization and implicit confidence calibration warrants deeper investigation. Our findings suggest that EM enhances model confidence by reinforcing high-probability reasoning paths 3.3. This implies EM might serve as a lightweight alternative to complex calibration techniques, especially for tasks where interpretability and robustness are critical. Developing evaluation protocols to more precisely quantify EM’s calibration effects will be an important direction.
@@
Entropy minimization is conceptually orthogonal to most existing post-training paradigms, including supervised fine-tuning (SFT) and reinforcement learning (RL). This opens exciting opportunities for hybrid methods. For example, EM can be applied before SFT to sharpen a model’s predictive distributions, improving its receptiveness to downstream supervision. Alternatively, EM could serve as a regularization strategy during SFT or RLHF.
@@ 5.1 Entropy Minimization
Agarwal et al.[20] was the first to discover that simply optimizing the entropy loss can lead to significant improvements in reasoning performance. However, they did not explore this phenomenon in depth, stating in the original paper, "We leave the rigorous analysis to future works." Wang et al.[1] is the first work to study entropy minimization in post-training for large language models. However, they consider the effectiveness of entropy minimization to be unreasonable, presenting experimental results but offering limited in-depth analysis of entropy minimization.
@@ 7 Acknowledgement
We thank Tian Xie, Zhengmao Ye, Peihao Wu and Derek Li at Ubiquant for their support and insightful discussions.